Final Program

Nikolaus Conference 2000 Darstellungstheorie-Tage 2000

(Lehrstuhl D für Mathematik, RWTH Aachen)

Thursday, December 7, 2000 2.01–2.32pm Thomas Breuer (Aachen): Computing Character Tables in the 21st Century (R220)

2.40–3.10pm Wolfgang Lempken (Essen): Linear Groups with a Special Class of Fixed Point Free Elements (R220)

coffee and discussion

4.30–5.06pm Marcos Soriano (Stuttgart): A Theorem on Extension Groups of Specht Modules for Type A Hecke Orders (Be225)

5.10–5.39pm  Jean Michel (Paris): Symmetric Braids (Be225)
7.00pm–?^{1}  Research Cambridge Style! (Dinner in “Agora”, Greek kitchen)

Friday, December 8, 2000 (morning talks in EA, afternoon talks AS, except 3.15pm)

10.05–10.35am Lacri Iancu (Lyon): Markov Traces and Generic Degrees in Type B_{n}
10.40–11.03am Joachim Simon (Stuttgart): On a New Proof of Tits’ Deformation Theorem
11.18–11.54am Gunter Malle (Kassel): Irreducibility of Symmetric and Alternating Squares

lunch break

2.15–2.47pm Martin Hertweck (Stuttgart): Units of p-Power Order in Principal Blocks of p-Constrained Groups

coffee break (Graduiertenkolleg)

3.15–4.15pm Meinolf Geck (Lyon): An Application of GAP to Algebraic Groups over Finite Fields (talk in the ‘Graduiertenkolleg’ seminar, Hörsaal IV)

4.51–5.20pm Sungsoon Kim (Paris): On the Blocks of the Cyclotomic Hecke Algebras of the Complex Reflection Group G(d, 1, r)

5.31–6.08pm Frank Lübeck (Aachen): Small Degree Representations of Finite Chevalley Groups in Defining Characteristic

7.00pm–?^{1}  Research Cambridge Style! (in the “Labyrinth”)

Saturday, December 9, 2000 (all talks in AS) 10.03–10.32am Harald Weber (Stuttgart): Structure of the Integral Group Ring for a Series of p-Groups

10.40–11.10am Bettina Eick (Kassel): The Orbit-Stabilizer Problem for Polycyclic Groups

11.20–11.45am Stefan Kohl (Stuttgart): Number of Orbits of Simple Groups under Automorphisms

lunch break

2.02–2.30pm  Burkhard Kuelshammer (Jena): Blocks with Abelian Defect Groups
2.40–3.18pm  Götz Pfeiffer (Galway): Implementing Character Tables of Reflection Groups
3.24–3.55pm  Rob Wilson (Birmingham): New Computations in the Monster

coffee, discussion, … — open end

Anyone was welcome to attend the talks. Frank Lübeck

1 Come next year to find out yourself!

final version (Mon Dec 11 10:40:06 CET 2000)